How do you solve y=6-2x
5y+8x=20

Answer by mananth(16946)

y=6-2x
2x+y=6
5y+8x=20
8x+5y=20
...
2x+1y=6 .............1
8x+5y=20.............2

multiply (1)by -4

-8x-4y =-24
8x+5y=20

Add the two equations

y=-4
plug value of y in 1
we get
2x-4=6
2x=10
x=5